OpenStudy (anonymous):

After lunch, two women drive in opposite directions. Moya drives east at 76 mi/hr. Indira drives west at 72 mi/hr. How far apart are they after 32 minutes?

OpenStudy (anonymous):

equations moya equation =76x indira -72x subtract two equation 76x-(-72x)=

OpenStudy (anonymous):

148?

OpenStudy (anonymous):

x represent time in hours so x=32/60 plug that in 76x-(-72x)= 76x+72x= x(76+72)= (32/60)(76+72)

OpenStudy (anonymous):

78.9333333

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok thanks